OpenStudy (anonymous):

How to solve: log(x+2) + log(x-1) = 1

OpenStudy (anonymous):

Hint: loga + logb = log(a/b)

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

Hmm give me a sec

OpenStudy (anonymous):

Not really :$ the answer is 3 and that just gives me a quadratic formula

OpenStudy (anonymous):

Are they base 10?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I assume

OpenStudy (anonymous):

1=log10

OpenStudy (anonymous):

So, log(x+2) + log(x-1) = 1 log{(x+2)(x-1)} = log10

OpenStudy (anonymous):

That may help

OpenStudy (anonymous):

Do you know what to do when both 2 sides have the same base?

OpenStudy (anonymous):

I don't think so (chlorophyll)

OpenStudy (anonymous):

So learn the rule, both logs will be CANCELLED out :)

OpenStudy (anonymous):

really? okay, but then I'm still confused :$

OpenStudy (anonymous):

okay, so it would still be (x+2)(x-1) = 10

OpenStudy (anonymous):

x^2 +x -2 = 10

OpenStudy (anonymous):

or x^2+x-12 actually

OpenStudy (anonymous):

=0

OpenStudy (anonymous):

... solve the quadratic equation?

OpenStudy (anonymous):

and then is the positive root just my answer?

OpenStudy (anonymous):

Perfect! I'm waiting for your x = ...?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Wow, why's so fast! Hmm...

OpenStudy (anonymous):

Anyway, correct!

OpenStudy (anonymous):

Questions?

OpenStudy (anonymous):

um no thanks :) will try that with the next question now!